OpenStudy (anonymous):

x 2 =x2 plz slove this question (x=2) but how??

OpenStudy (anonymous):

Is it \[x^2 = 2x \]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then? Can you type it in the TeX format,please?

myininaya (myininaya):

aruun you never answered my question. was it what I typed on the earlier thread?

OpenStudy (anonymous):

ok

myininaya (myininaya):

\[2^x=x^2\] was my interpretation

myininaya (myininaya):

please say if this is right or wrong?

OpenStudy (anonymous):

yup this right

myininaya (myininaya):

well what is the only number where the bases are the same and where the exponents are the same?

myininaya (myininaya):

 right?

myininaya (myininaya):